OpenStudy (anonymous):

Given the points (2, 3) and (6, 8), write the point-slope form of the equation. A. y - 3 = x - 2 B. y - 3 = 5/4 (x - 2) C. y - 8 = 4/5 (x - 6)

ganeshie8 (ganeshie8):

start wid finding the slope

OpenStudy (anonymous):

-6.5

ganeshie8 (ganeshie8):

doesnt look correct

ganeshie8 (ganeshie8):

slope = (difference in y) / (difference in x)

ganeshie8 (ganeshie8):

= (8-3)/(6-2) = 5/4

OpenStudy (anonymous):

Okay.

ganeshie8 (ganeshie8):

thats the slope, do u knw the equation of line in point-slope form ?

OpenStudy (anonymous):

B

ganeshie8 (ganeshie8):

good guess :)